[Brahms-l] Shift report 20050418 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Apr 18 2005 - 08:02:19 EDT
Shifter leader was SB
Shifters were 


1. Physics running with beam from previous shift continued 
until dump at 2:05 am.  Physics running since 3:20 with a 
50 minute machine set up time in between while beam was on.

2. 5.5 hours of physics running.
Runs 14073,14074,14075,14076,14077

3. No detector problems


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 18 08:02:31 2005

This archive was generated by hypermail 2.1.8 : Mon Apr 18 2005 - 08:02:44 EDT